If ๐‘“ is a continuous function from a metric space X into a metric space Y and {๐‘ฅ๐‘›} is a sequence in X

which converges to ๐‘ฅ then show that the sequence {๐‘“(๐‘ฅ๐‘›)} converges to ๐‘“(๐‘ฅ) in Y.


1
Expert's answer
2021-07-26T08:01:19-0400

๐‘“ is a continuous function from a metric space X into a metric space Y if and only if "f" is continuous at every point of X.


"f(x)" is the limit of the sequence "f(x_n)" if and only if

(1) "\forall \varepsilon>0\, \exists N\in\mathbb{N}\, \forall n>N\, d_Y(f(x),f(x_n))<\varepsilon"


Fix "\varepsilon>0". Since ๐‘“ is a continuous function at point "x\in X" then

(2) "\exists\delta>0\, \forall x': d_X(x,x')<\delta\to d_Y(f(x),f(x'))<\varepsilon"


Since "x=\lim\limits_{n\to\infty}x_n" , there exists "N\in\N" such that for all "n>N""d_X(x,x')<\delta".

Applying (2), we get "d_Y(f(x),f(x_n))<\varepsilon".


Finally, we have deduced (1), and the assertion is proved.